What applied horizontal force is required to accelerate a 5 π‘˜π‘” along a horizontal surface. With an acceleration of 2 π‘š/𝑠 2 if the coefficient of friction is 0.15.


1
Expert's answer
2021-10-08T09:28:25-0400
"F_{appl}-F_{fr}=ma,""F_{appl}-\mu mg=ma,""F_{appl}=m(\mu g+a),""F_{appl}=5\ kg\cdot(0.15\cdot9.8\ \dfrac{m}{s^2}+2.0\ \dfrac{m}{s^2})=17.35\ N."